The area of a rectangle is 55.96 meters square. If the length is 16 meters, what is its perimeter?
Answer by TimothyLamb(4379)

Lw = 55.96
w = 55.96/L
L = 16
---
w = 55.96/16
w = 3.4975
---
p = 2L + 2w
p = 2*16 + 2*3.4975
p = 38.995 meters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php